# Recent Advances in OOD Detection: Problems and Approaches

### v2

@@ 6.3.2. Natural Language Processing
OOD detection is also explored in various tasks across numerous Natural Language Processing applications. The two most common applications are as follows:
  • Intent Detection. Intent Detection is a significant application of OOD detection in NLP. The datasets used for evaluation include CLINC150  (Larson et al., 2019), Banking  (Casanueva et al., 2020), StackOverflow  (Xu et al., 2015), among others.
  • Text Catagory. In text category OOD detection applications, datasets like News Category  (Misra, 2022) and SST-2 (Socher et al., 2013) are commonly utilized to form ID/OOD pairs and assess the models’ detection capabilities.
@@ 6.3.3. Beyond Computer Vision and Natural Language Processing
  • Audio data. In audio OOD detection, MSCW (Micro-EN) (Mazumder et al., 2021) and Vocalsound  (Gong et al., 2022) are usually used as ID datasets, and they also act as OOD for each other.
  • Graph data. Recent studies have proposed various approaches for OOD detection in graph data. The existing graph-level OOD detection benchmark comprises a total of 10 pairs of datasets from TU  (Morris et al., 2020) and OGB datasets (Hu et al., 2020), which has been widely utilized.
  • Reinforcement learning. Currently, there is a rising trend of integrating OOD detection with reinforcement learning  (Leonardos et al., 2023) to bolster model robustness.  Mohammed and Valdenegro-Toro (2021) provide a benchmark and explore methods for crafting tailored reinforcement learning environments that can generate OOD data.
@@ 7.1. Better methodologies of OOD detection
Meta-Learning Adaptation. Faced with the quick adaptation challenge in test-time OOD detection, mete-learning algorithms, which provide a learning-to-learn paradigm to efficiently adapt the model to new test data, may be the solution. Additionally, in addressing the dauntingly vast sample space of potential outliers inherent in training-driven OOD detection, improved sampling methods could offer a pathway to the efficient utilization of outliers  (Ming et al., 2022a).
Theoretically-Driven Score Designing. In the traditional single-modal regime, numerous post-hoc scores have been carefully designed for OOD detection, effectively reducing training costs. However, as the field has progressed into the multi-modal domain, there is an increasing demand for theoretically-driven score designs. MCM  (Ming et al., 2022b) is a notable example, but it is merely an extension of Softmax and does not deeply explore the relationships between text and image. Therefore, more advanced scores are needed.
@@ 7.2. More practical scenarios of OOD detection
Quick Test-Time Adaptation. The advent of test-time scenarios like CAOOD  (Wu et al., 2023) significantly advances the application of OOD detection in real-world contexts, promising improved reliability and adaptability.
Multi-Modal Detection. Exploring multi-modal OOD detection enhances our grasp of data dynamics and boosts model efficacy across varied sensory inputs. Additionally, the forthcoming integration of LLMs into multi-modal OOD detection stands to transform the field, marrying advanced linguistic analysis with other data types to forge more effective and versatile detection systems for complex applications.
Open-Vocabulary Scenario. Existing LPM-based methods assume that ID categories are accessible. However, this assumption does not hold in open-vocabulary scenarios.  Li et al. (2024b) addresses this by learning migratable negative prompts. It uses a subset of ID labels for prompt learning, and these learned prompts can be applied to other unseen labels. OOD detection in open-vocabulary settings further relaxes the requirement for prior access to ID knowledge, paving the way for the development of more generalized models.
Noise Settings. While previous research has largely focused on standard clean label settings, few studies have explored the other side of the coin: noisy label scenarios.  Humblot-Renaux et al. (2024) investigate practical scenarios and examine the impact of varying levels of noise on OOD detection. Then they provide several key takeaways for improving OOD detection in noisy label environments.
@@ 7.3. New applications of OOD detection
Additional Modalities. While OOD detection has made strides across diverse sectors, its potential in speech and physiological signal analysis remains largely untapped. In particular, OOD issues are prevalent in emotion-related physiological signals due to the variability in subjects’ emotional states. Therefore, this represents a promising area for further exploration. Some of the current methods  (Dong et al., 2024) also provide benchmarks for reference.
Human-in-the-loop Application. A pivotal future direction, proposed by  Vishwakarma et al. (2024), is the integration of human insight into the detection process. This human-in-the-loop approach, particularly in high-stakes decision-making, will be vital for enhancing the accuracy and responsiveness of OOD detection systems, merging human intuition with algorithmic precision.
Web Image Scraping. To automate the process of scraping images from the Internet,  Miyai et al. (2023c) propose a zero-shot ID detection task, a concept derived from zero-shot OOD detection. An image will be classified as an ID image if it contains ID objects; it will only be considered an OOD image if it lacks any ID objects. It presents a novel perspective on the application of OOD detection and merits further exploration
@@ 8. Conclusion
OOD detection is critical for trustworthy machine learning. In this paper, we provide a comprehensive review of recent advances in OOD detection, focusing for the first time on the problem scenario perspective: training-driven, training-agnostic, and large pre-trained model-based OOD detection. We also summarize extensively used evaluation metrics, experimental protocols, and diverse applications. We believe that our novel taxonomy of existing papers and extensive discussion of emerging trends will contribute to a better understanding of the current state of research, assist practitioners in selecting suitable approaches, and inspire new research hotspots.

### v3

@@ 6.3.2. Natural Language Processing
OOD detection has become a crucial research topic in NLP, especially with the advent of LLMs and their deployment in real-world applications (Liu et al., 2025c; Lang et al., 2023). Below, we summarize key tasks and representative methods:
  • Intent Detection. Intent Detection is a significant application of OOD detection in NLP, particularly in dialogue systems, where identifying user intentions that fall outside the predefined set of intents is crucial (Wang et al., 2024c; Zhan et al., 2022; Akbari et al., 2025). Common datasets for evaluation include CLINC150 (Larson et al., 2019), Banking77 (Casanueva et al., 2020), and HWU64 (Liu et al., 2020b). These datasets are specifically designed to evaluate OOD detection in intent classification tasks.
  • Text Classification. In text classification tasks, OOD detection helps identify texts that do not belong to any of the known categories. Datasets such as News Category (Misra, 2022) and SST-2 (Socher et al., 2013) are frequently used for in-distribution data, with 20 Newsgroups (20NG) (Lang, 1995) often serving as an OOD dataset. Methods like VI-OOD (Zhan et al., 2024) enhance detection by optimizing joint distributions.
  • Question-Answering Systems. Question-answering systems require robust mechanisms to handle diverse user queries while maintaining response accuracy. The challenge of identifying questions that fall outside the system’s knowledge domain has led to sophisticated OOD detection approaches. Using standard in-distribution datasets (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), PubMedQA (Jin et al., 2019)) and OOD datasets (SQUAD 2.0 (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2019)), recent LLM-based methods (Zhang et al., 2024c) leverage likelihood ratios between pretrained and fine-tuned models for effective OOD detection.
  • Mathematical Reasoning. Mathematical problem-solving systems face challenges when encountering problem types not seen during training. Using MultiArith (Roy et al., 2016) as the primary in-distribution dataset, researchers evaluate OOD detection against diverse benchmarks including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), SVAMP (Patil et al., 2021), AddSub (Hill et al., 2015), SingleEq (Wang et al., 2024b), and SingleOp (Wang et al., 2024b). Recent advances, such as TV-Score (Wang et al., 2024b), analyze embedding trajectory variations to enhance OOD detection performance.
  • Document Classification. In (multimodal) document classification, where documents integrate both text and images, OOD detection is essential for identifying unseen document types. Datasets such as Tobacco3482 (Aggarwal et al., 2017) and FinanceDocs (Constantinou et al., 2024) are commonly used for evaluation. The AHM (Constantinou et al., 2024) approach enhances OOD detection by improving the representation of multimodal features.
@@ 6.3.3. Beyond Computer Vision and Natural Language Processing
  • Audio Recognition. In the field of audio recognition, OOD detection is essential for identifying inputs such as synthesized speech and dialects that diverge from the training distribution. Traditional probabilistic scoring methods often underperform on atypical audio samples (Lane et al., 2006), motivating the adoption of alternative approaches including confidence-based topic classification (Lane et al., 2006), neural sentence embeddings (Ryu et al., 2017), and deep nearest neighbor analysis (Bukhsh and Saeed, 2023). Autoencoder architectures built upon WavLM (Du et al., 2024b) have demonstrated improved performance, particularly in detecting synthetic speech. Standard datasets for evaluating OOD detection in this domain include MSCW (Mazumder et al., 2021) and Vocalsound (Gong et al., 2022), which provide both in-distribution and out-of-distribution samples for comprehensive benchmarking.
  • Graph Neural Networks. Within graph neural networks (GNNs), OOD detection encompasses both node-level and graph-level tasks. Commonly explored techniques include uncertainty-based graph classification (Bazhenov et al., 2022), energy-based detection (Wu et al., 2023b; Jiang et al., 2025), and semi-supervised learning (Huang et al., 2022b). The HGOE framework (Junwei et al., 2024) enhances graph-level detection by incorporating anomalous samples during training, while PGR-MOOD (Shen et al., 2024) utilizes diffusion models to identify out-of-distribution molecular structures. GOLD (Wang et al., 2025b) generates pseudo-OOD data with significant differences from the ID embeddings by using a latent generative model to mimic the embeddings of training data and training it jointly with a GNN encoder and an OOD detector.Benchmarking typically relies on paired datasets from TU (Morris et al., 2020) and OGB (Hu et al., 2020), which facilitate systematic evaluation of OOD detection methods in GNNs.
  • Reinforcement Learning. OOD detection plays a crucial role in reinforcement learning (RL) by enhancing system robustness when encountering unprecedented states or actions (Leonardos et al., 2023). Common approaches include uncertainty quantification through dropout techniques and bootstrapping methods (Sedlmeier et al., 2019), as well as systematic environment modifications (Mohammed and Valdenegro-Toro, 2021). In the context of offline RL, advanced methods such as ADAC (Chen et al., 2025b) systematically assess OOD actions to effectively address distribution shifts. Comprehensive evaluation frameworks for OOD detection in RL primarily concentrate on various forms of environmental perturbations (Mohammed and Valdenegro-Toro, 2021), providing standardized testing scenarios for robustness assessment.
  • Time Series Analysis. In the domain of time series analysis, the inherent temporal dependencies present significant challenges for conventional OOD detection approaches. SRS (Belkhouja et al., 2023) addresses this by leveraging time series decomposition in conjunction with deep generative models to estimate conditional likelihoods, while CODiT (Kaur et al., 2023) introduces temporal equivariance to better capture sequential patterns. Recent studies have further evaluated modality-agnostic algorithms on multivariate time series datasets, outlining persistent challenges and future research directions (Gungor et al., 2025). In practice, OOD detection has been widely adopted in fraud detection services, such as the AWS Fraud Detector (web, ), which utilizes machine learning to identify transaction patterns that deviate from historical norms, effectively flagging potentially fraudulent activities in real time.
  • Biological Classification. In biological classification, OOD detection plays a vital role in identifying previously unseen species or anomalies that extend beyond the categories present in training data. The integration of DNA barcodes has been shown to enhance image-based classification models through re-ranking strategies, thereby improving robustness and detection accuracy (Arslan et al., 2024).
  • Spiking Neural Networks. Spiking Neural Networks (SNNs), recognized for their event-driven efficiency, present distinctive challenges for OOD detection due to their sparse latent representations. Recent methods have proposed combining feedforward learning with distance-based scoring and gradient-free attribution to address these challenges (Terres-Escudero et al., ). The N-MNIST dataset (Orchard et al., 2015) is commonly used for evaluation, with OOD samples generated through diverse event patterns.
  • Astronomical Imaging. In astronomical imaging, OOD detection is essential for distinguishing between simulated and real galaxy images, where conventional statistical techniques often prove inadequate. Approaches such as Bayesian model comparison and sparse variational autoencoders (Zhou et al., 2024) have been utilized to uncover subtle discrepancies. The SDSS dataset is widely adopted for benchmarking, with simulation-induced biases serving as OOD samples.
  • Edge Computing. Within edge computing, efficient and privacy-preserving OOD detection is critical for devices with limited computational resources. SecDOOD (Li et al., 2025) addresses these constraints by employing a collaborative cloud-device architecture and HyperNetwork-based parameter generation, thereby eliminating the need for device-side backpropagation. CIFAR-10 and Tiny ImageNet are standard datasets for in-distribution and OOD evaluation in this context.
@@ 7.1. Better methodologies of OOD detection
Meta-Learning Adaptation. Rapid adaptation to novel data distributions at test time remains a significant challenge in OOD detection. Meta-learning algorithms, which follow a learning-to-learn paradigm, have emerged as promising solutions for enabling efficient adaptation to unseen test data. Additionally, the vast sample space of potential outliers in training-driven OOD detection motivates the development of improved sampling strategies for more effective utilization of outlier information (Ming et al., 2022a).
Model Selection. Selecting the most suitable OOD detection approach for a given task is non-trivial, as detection performance often varies across data distributions. Automated model selection frameworks, such as MetaOOD (Qin et al., 2025), leverage meta-learning to facilitate zero-shot model selection, while DSDE (Geng et al., 2024) and score-combining methods (Reehorst and Schniter, 2025) utilize proportion estimation and likelihood ratio tests to enhance ensemble and score fusion strategies. Interestingly, recent work has shown that partially trained, immature models can exhibit effective OOD detection capabilities (Montazeran and Köthe, 2025), suggesting new directions for leveraging model dynamics in detection.
Active Learning. Active learning strategies, which prioritize labeling the most informative data points, offer efficiency gains in data-scarce settings. In OOD detection, active learning can focus annotation efforts on ambiguous or high-uncertainty samples, thereby improving the model’s discrimination between in-distribution and out-of-distribution data. SISOM (Schmidt et al., 2024) exemplifies a unified approach that combines active learning and OOD detection by selecting samples based on feature space distances.
Continual and Incremental Learning. As real-world data is often encountered in a streaming and evolving manner, continual and incremental learning paradigms are increasingly relevant for OOD detection. These approaches enable models to adapt to new tasks and data distributions over time while retaining knowledge of previous ones. Recent efforts include continual unsupervised OOD detection (Doorenbos et al., 2024), systematic evaluation benchmarks like OpenCIL (Miao et al., 2024b), and hierarchical two-sample tests for continual OOD detection (Liu et al., 2025d).
Theoretically-Driven Score Design. While many post-hoc scoring methods have been proposed for OOD detection in single-modal regimes, the transition to multi-modal domains necessitates more principled, theoretically motivated score designs. For example, MCM (Ming et al., 2022b) extends softmax-based approaches, but further innovation is needed to better capture the complex relationships in multi-modal data.
@@ 7.2. More practical scenarios of OOD detection
Quick Test-Time Adaptation. The introduction of test-time adaptation scenarios, such as CAOOD (Wu et al., 2023a), marks a step forward in making OOD detection more reliable and adaptable in dynamic real-world environments.
Federated Environments. In practical scenarios, data is often distributed across multiple clients with heterogeneous distributions. Federated learning (FL) enables collaborative model training while preserving privacy, but also brings unique OOD detection challenges due to covariate and semantic shifts among clients. Addressing these challenges is crucial for robust federated models. For instance, FOOGD (Liao et al., 2024) combines score matching and feature invariance to handle simultaneous OOD shifts in federated settings.
Multi-Modal Detection. The proliferation of multi-modal data—ranging from images and text to audio and video—demands OOD detection methods capable of handling diverse inputs. Benchmarks such as MultiOOD (Dong et al., 2024) provide comprehensive testbeds for multi-modal OOD detection, introducing algorithms like A2D and NP-Mix for challenging scenarios. Further advancements, including DIAEF (Gao et al., 2024) and DPU (Li et al., 2024d), focus on cross-modal alignment and dynamic prototype updating, respectively. The integration of large language models (LLMs) into multi-modal OOD detection is poised to further enhance the field by combining advanced linguistic reasoning with other modalities.
Multi-Label OOD Detection. Real-world applications often involve multi-label data, where each instance may belong to multiple categories. Detecting OOD samples in these scenarios is challenging due to increased dimensionality and label correlations. Recent approaches (Mei et al., 2024; Aguilar et al., 2025) leverage uncertainty modeling and subjective logic to improve robustness in multi-label OOD detection.
Open-Vocabulary Scenario. Traditional approaches often assume access to all in-distribution categories, an assumption that does not hold in open-vocabulary contexts. Recent work (Li et al., 2024c) explores learning migratable negative prompts, enabling OOD detection even when only a subset of in-distribution labels is available, thus paving the way for more generalizable models.
Noisy Label Environments. Most prior studies focus on clean label settings, but real-world data often contains label noise. Investigations into noisy label scenarios (Humblot-Renaux et al., 2024) offer insights and practical recommendations for improving OOD detection robustness under such conditions.
Long-Sequence Processing. In NLP, the detection of OOD inputs becomes increasingly complex when language models process sequences that exceed their pre-trained context window. In such scenarios, models frequently encounter unfamiliar dependency structures, increased attention demands, and altered positional encoding behaviors, all of which deviate from patterns seen during training. These distributional shifts impede the model’s generalization capabilities, highlighting the importance of thoroughly understanding their effects. Advancing OOD detection in this context is essential for ensuring reliable model performance on sequences that extend beyond standard context boundaries (Liu et al., 2025c).
@@ 7.3. New applications of OOD detection
Additional Modalities. While significant advances have been made in vision and language domains, OOD detection in modalities such as speech and physiological signals remains underexplored. These areas, particularly emotion-related physiological signals with high inter-subject variability, present promising directions for future research (Dong et al., 2024).
Human-in-the-Loop Systems. The integration of human expertise into OOD detection is emerging as a critical direction, especially in high-stakes or safety-critical applications. Recent frameworks (Vishwakarma et al., 2024; Bai et al., 2024b, c) incorporate human feedback to refine detection thresholds and guide annotation, balancing algorithmic precision with human intuition. The AHA framework, for instance, strategically targets ambiguous regions for selective annotation, improving generalization and error correction under limited supervision.
Web Image Scraping. OOD detection is also finding novel applications in automating web image scraping. For example, zero-shot ID detection (Miyai et al., 2023c) classifies images as in-distribution if they contain relevant objects, offering a fresh perspective on leveraging OOD detection for large-scale, real-world data acquisition.
@@ 8. Conclusion
OOD detection is a critical component for trustworthy machine learning. In this paper, we provide a comprehensive review of recent advances in OOD detection, focusing for the first time on the problem scenario perspective: training-driven, training-agnostic, and large pre-trained model-based OOD detection. We also summarize extensively used evaluation metrics, experimental protocols, and diverse applications. We believe that our taxonomy of existing papers and extensive discussion of emerging trends will contribute to a better understanding of the current state of research, assist practitioners in selecting suitable approaches, and inspire new research hotspots.
9. Acknowledgements
We sincerely appreciate the constructive comments and suggestions provided by all the reviewers and the editor, which have greatly improved the quality of our work. We are truly grateful for their valuable time and insightful feedback.